Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution.
A sum of money amounting to $3.80 consists of dimes and quarters. If there are 20 coins in all, how many are quarters?

Answer by mananth(16946)

Let the number of dimes be x
the number of quarters be y
x+y =20
10x+25y=380 cents
10x+10y -(10x+25y)=200 -380
10x+10y-10x -25y=-180
-15y= -180
y=-180/-15
y=12 the number of quarters
20-12 = 8 will be the number of dimes